The electric potential due to a charged sphere of radius 2.0 cm is 400V at a distance of 10 cm from the centre. What will be the electric field at a distance of 20 cm from the centre of the sphere?

Grade:12th pass

2 Answers

parthasarathy
9 years ago
V=E*D
E=V/D
E=400/0.2=2000 N/m
 
parthasarathy
9 years ago
V=E*D
E=V/D
E=400/0.2=2000 N/m